OpenStudy (anonymous):

Really simple question. 16x^2 = y. What is x? what is y?

OpenStudy (luigi0210):

well y is given x is \[-/+\frac{ \sqrt{y} }{ 4 }\]

OpenStudy (anonymous):

Thank you. but what is the answer?

OpenStudy (luigi0210):

y=16x^2 x=(+/-sqrt y)/4`